A bridge is designed to hold a certain load. If we increase all the bridge's dimensions by a factor of 5, will the bridge hold the mass that it was originally intended to hold? Explain. A box weighs 81 more times than it did before I enlarged it. How much did I enlarge the box?
1
Expert's answer
2019-12-06T09:40:11-0500

1) Yes. It can be false if we decrease the dimensions.


2)


"\frac{V'}{V}=\frac{m'}{m}=\left(\frac{a'}{a}\right)^3=81"

"\frac{a'}{a}\approx4.33"